# Deep polytopic autoencoders for low-dimensional linear parameter-varying approximations and nonlinear feedback design

### v1

@@ 6 Numerical experiments
Figure 1: Grid search for the selection of $r$ and $q$ with lowest averaged reconstruction error.
Table 1: Model complexity of the different PAE schemes and POD: # stands for number of. Further, we distinguish between linear (lin.) and nonlinear (nonl.) mappings in the en–/decoding.
@@
We consider the same problem, numerical setup, and the same data points for training the models as in [16]. In short, the flow control problem of stabilizing the unstable steady-state of the flow behind a cylinder is modeled through the difference $v$ of the velocity to the steady-stated and observed through outputs $y=Cv$. It is determined by the semi-discrete Navier-Stokes equations
$M\dot{v}=N(v)v+Bu,\quad y=Cv,\quad v(0)=0\in\mathbb{R}^{n},$ (15)
[…]
Figure 2: Snapshots of the flow simulation setup.
We consider the steady-state solution as target state and aim for the stabilization of the difference system eq. 15. In order to trigger instabilities and for the collection of training data, a test input
@@
is applied. As described in [14], the chosen deep convolutional encoder leverages depthwise separable convolutions along with a sparse interpolation matrix to mitigate computational complexity and memory requirements in this rather high-dimensional setup. As a result the model complexity in terms of parameters that define the encoder and decoder can be lower than for, say, a comparable POD reduction; see Table 1. For training the PAEs, we adopt a semi-supervised learning methodology with a combined loss function
$\mathcal{L}(v^{(i)}):=\lambda\|\tilde{v}^{(i)}-v^{(i)}\|_{M}-l(\rho^{(i)})\cdot\log(c(\rho^{(i)})),$
where $v^{(i)}$ is the $i$-th data point, $\rho^{(i)}:=\mu(v^{(i)})$ and $\tilde{v}^{(i)}:=\nu(\rho^{(i)})$ for the corresponding encoder and decoder $\mu$ and $\nu$, where $l$ assigns labels that were precomputed by $k$-means clustering in $\mathbb{R}^{r}$, and where $\lambda>0$ is a weighting factor between the reconstruction error (measured in the $M$-weighted discrete $L^{2}$-norm) and the clustering error. We note that the chosen cross-entropy loss is commonly used to smoothly assign labels and that the labeling $l$ is defined such that the target $\rho^{*}=0$ is assigned to the first cluster. As it is standard practice, the loss is computed as the average over a batch of data points. In our experiments, we consider $401$ data points equally distributed in the time interval $[0,0.5]$ as training data and we use the Adam optimizer with a learning rate of $0.005$, with $3200$ epochs, a batch size of $64$ and the loss weight $\lambda=100$.
@@ 6.1 Reconstruction errors
$\frac{1}{N}\sum_{i=1}^{N}\|v^{(i)}-\tilde{v}^{(i)}\|_{M},$
where $v^{(i)}$ is the $i$-th data point and $\tilde{v}^{(i)}=\nu(\mu(v^{(i)}))$ for the corresponding encoder and decoder $\mu$ and $\nu$. As laid out in Table 2, the PAE-1.r is comparable to the standard linear approach of POD for low dimensions $r=2,3,4$, whereas the nonlinear encoding through the cluster selection significantly improves the reconstruction in low dimensions. As for the series expansion of the control law, only the first-order approximation is relevant, we tabulate this reconstruction error as well. Note that the bad average performance is due to the strong deviation for larger $t$, i.e. larger values of $\rho$, whereas for small $t$, the first-order approximation closely aligns with the nonlinear approach; see Figure 3.
Figure 3: Trajectory with pointwise reconstruction errors $\|v(t^{(i)})-\tilde{v}(t^{(i)})\|_{M}$. The improvement of using $3$ clusters is clearly visible as is the initial alignment with the nonlinear approach PAE-3.5 (and the subsequent deviation from it) of the first-order expansion.
@@ 6.2 SDRE approximation through series expansions
We check the approximation of the SDRE feedback along the example trajectory $v$ generated by the test input eq. 16. For that we compute the true SDRE feedback $u_{P}(t)=-B^{\mkern-1.5mu\mathsf{T}}P(v(t))v(t)$, where $P$ solves the SDRE, and compare to $u_{\tilde{P}}$ and $u_{P_{|\alpha|\leq p}}$ that denote the feedback computed through the LPV approximation in eq. 9 (via PAE-q.5, $q=1,3$) and the expansions eq. 14 for $p=0,1,2$, respectively. The data in Figure 4 shows that the LPV approximation works particularly well for $t\leq 0.5$, which is the training regime, and that the first and second-order expansions significantly improve the zero-order approximation, which would be the LQR gain, but with a slight effect of the additional second-order terms.
@@
Figure 4: Differences to the exact SDRE feedback along a sample trajectory for feedback approximations through the LPV approximation of the nonlinearity by PAE-q.5, for $q=1,3$, and the corresponding zeroth, first and second-order series expansions of the SDRE feedback.
@@ 6.3 Feedback performance
$\frac{1}{t_{e}}\Big{(}\int_{t_{s}}^{t_{e}}\|B^{\mkern-1.5mu\mathsf{T}}\bigl{(}\sum_{|\alpha|\leq p}\rho(v(t))^{(\alpha)}P_{\alpha}\bigr{)}v(t)\|^{2}\operatorname{d}t\Big{)}^{\frac{1}{2}},$ (17)
that tends to 00 for $t_{e}\to\infty$ if stabilization is achieved, and that is finite if and only if the simulation has not blown up before a finite end time $t_{e}$.
@@
Figure 6: Measured output and feedback input signals for the distinguished data points from Figure 5.
We plot the values of the performance index eq. 17 for $t_{e}=7.5$ and for various approximations to the nonlinear SDRE feedback in Figure 5 and the trajectories of selected cases in Figure 6. As we have observed before, the $p=0$ expansion in eq. 14, i.e., the LQR feedback, is remarkably performant for the optimal choice of the Tikhonov parameter $\gamma$. However, considering higher order expansions with $p=1,2$ together with the option to adapt the underlying approximation scheme, significantly widens the domain of attraction and can ensure reliable performance in particular for larger values of $\gamma$; see, e.g., Figure 6 (top row) for the trajectory data for a setup with $\gamma=10$, where only the second-order expansion achieves stabilization.
@@ 7 Conclusion
In the forward simulation model, the proposed nonlinear autoencoder with smooth clustering significantly improves the parametrization quality at very low dimensions. For the controller design, it was established that for second-order expansions of the SDRE feedback law, only the first-order expansions of the reduced-order LPV coefficient are relevant, which reduces the computational effort by large. Concerning the performance of the resulting feedback, the overall picture is diffuse. While the linear LQR approach, for suitable weight parameters, ensured stability also well away from the origin, the nonlinear feedback design based on different approximation schemes, clearly extended the basin of attraction. As the choice of the approximation was solely based on a single open-loop simulation in a time range that did not cover the range considered in the closed-loop simulation, in future work, we expect to gain additional performance by optimizing the approximation in view of feedback control.

### v2

@@ 6 Numerical experiments
6.1 Numerical setup
Figure 1: Spatial domain and snapshots of the flow behind a cylinder in the unstable steady state (left) and in the periodic vortex shedding regime (right); cf. [30].
For comparability reasons, we consider here the same problem, numerical setup, and the same data points for training the models as in [30]. In short, the flow control problem of stabilizing the unstable steady-state of the flow behind a cylinder is modeled through the difference of the velocity $v$ to the steady-stated and observed through outputs $y=Cx$. It is determined by the semi-discrete Navier-Stokes equations
$M\dot{x}=N(x)x+Bu,\quad y=Cx,\quad x(0)=0\in\mathbb{R}^{n},$ (23)
[…]
Table 1: Model complexity of the different PAE schemes and POD: # stands for number of. We distinguish between linear (lin.) and nonlinear (nonl.) mappings in the encoding and decoding.
@@ 6.1 Numerical setup
We consider the steady-state solution as target state and aim for the stabilization of the difference system eq. 23. To trigger instabilities and to collect training data, a test input
@@
is applied. We choose a deep convolutional encoder with depthwise separable convolutions to cope with the high computational complexity and memory requirements in this rather high-dimensional setup; see [32] for details. As a result the model complexity in terms of parameters that define the encoder and decoder can be lower than for, say, a comparable POD reduction; see Table 1. For training the PAEs, we adopt a semi-supervised learning methodology with the loss function
$\mathcal{L}(v^{(i)}):=\lambda\|\tilde{x}^{(i)}-x^{(i)}\|_{M}-\ell(\rho^{(i)})\cdot\log(c(\rho^{(i)})),$
where $x^{(i)}$ is the $i$-th data point, $\rho^{(i)}:=\mu(x^{(i)})$ and $\tilde{x}^{(i)}:=\nu(\rho^{(i)})$ for the corresponding encoder and decoder $\mu$ and $\nu$, where $\ell$ assigns labels that were precomputed by $k$-means clustering in $\mathbb{R}^{r}$, and where $\lambda>0$ is a weighting factor between the reconstruction error (measured in the $M$-weighted $L^{2}$-norm) and the clustering error. We note that the chosen cross-entropy loss is commonly used to smoothly assign labels and that the labeling $\ell$ is defined such that the target $\rho^{\ast}=0$ is assigned to the first cluster. As it is standard practice, the loss is computed as the average over a batch of data points. Here, we consider $401$ data points equally distributed in the time interval $[0,0.5]$ as training data, and we use the Adam optimizer with learning rate $0.005$, $3200$ epochs, a batch size of $64$ and the loss weight $\lambda=100$.
@@ 6.2 Reconstruction errors
Figure 2: Trajectory with pointwise reconstruction errors $\|x(t^{(i)})-\tilde{x}(t^{(i)})\|_{M}$. The improvement of using $3$ clusters is clearly visible as is the initial alignment with the nonlinear approach PAE-3.5 (and the subsequent deviation from it) of the first-order expansion.
@@
$\frac{1}{N}\sum_{i=1}^{N}\|x^{(i)}-\tilde{x}^{(i)}\|_{M},$
where $x^{(i)}$ is the $i$-th data point and $\tilde{x}^{(i)}=\nu(\mu(x^{(i)}))$ for the corresponding encoder and decoder $\mu$ and $\nu$. As laid out in Table 2, the PAE-1.r is comparable to the standard linear approach of POD for low dimensions $r=2,3,4$, whereas the nonlinear encoding through the cluster selection significantly improves the reconstruction in low dimensions. As for the series expansion of the control law, only the first-order approximation is relevant, we tabulate this reconstruction error as well. Note that the bad average performance is due to the strong deviation for larger $t$, i.e., larger values of $\rho$, whereas for small $t$, the first-order approximation closely aligns with the nonlinear approach; see Figure 2.
Figure 3: Grid search for the selection of $r$ and $q$ with lowest averaged reconstruction error. A trade-off between the number of clusters and the reduced dimension becomes clearly visible.
@@ 6.3 SDRE approximation through series expansions
Figure 4: Differences to the exact SDRE feedback along a sample trajectory for feedback approximations through the LPV approximation of the nonlinearity by PAE-q.5, for $q=1,3$, and the corresponding zeroth, first and second-order series expansions of the SDRE feedback. The LPV approximation provides the most accurate approximation to the true SDRE feedback in its training regime up to $0.5$. The first and second-order approximations also strongly improve over the classical LQR feedback.
We check the approximation of the SDRE feedback along the example trajectory $x$ generated by the test input eq. 24. For that we compute the true SDRE feedback
$u_{P}(t)=-B^{\mkern-1.5mu\mathsf{T}}P(x(t))x(t),$
where $P$ solves the SDRE, and compare to $u_{\tilde{P}}$ and $u_{P_{|\alpha|\leq p}}$ that denote the feedback computed through the LPV approximation in eq. 11 (via PAE-q.5, $q=1,3$) and the expansions eq. 17 for $p=0,1,2$, respectively. The data in Figure 4 shows that the LPV approximation works particularly well for $t\leq 0.5$, which is the training regime, and that the first and second-order expansions significantly improve the zero-order approximation, which would be the LQR gain, but with a slight effect of the additional second-order terms.
@@ 6.4 Feedback performance
Figure 6: Measured output and feedback input signals for the distinguished data points from Figure 5. The second-order LPV approximation is capable of stabilizing the nonlinear system in situations, in which the first-order approximation begins to fail.
$\frac{1}{t_{\operatorname{e}}}\Big{(}\int_{t_{s}}^{t_{\operatorname{e}}}\|B^{\mkern-1.5mu\mathsf{T}}\bigl{(}\sum_{|\alpha|\leq p}\rho(x(t))^{(\alpha)}P_{\alpha}\bigr{)}x(t)\|^{2}\operatorname{d}t\Big{)}^{\frac{1}{2}},$ (25)
that tends to 00 for $t_{\operatorname{e}}\to\infty$ if stabilization is achieved, and that is finite if and only if the simulation has not blown up before a finite end time $t_{\operatorname{e}}$.
We plot the values of the performance index eq. 25 for $t_{\operatorname{e}}=7.5$ and for various approximations to the nonlinear SDRE feedback in Figure 5 and the trajectories of selected cases in Figure 6. As we have observed before, the $p=0$ expansion in eq. 17, i.e., the LQR feedback, is remarkably performant for the optimal choice of the Tikhonov parameter $\gamma$. However, considering higher order expansions with $p=1,2$ together with the option to adapt the underlying approximation scheme, significantly widens the domain of attraction and can ensure reliable performance in particular for larger values of $\gamma$; see, e.g., Figure 6 (top row) for the trajectory data for a setup with $\gamma=10$, where only the second-order expansion achieves stabilization. The reason might be that a larger $\gamma$ seeks for small (in magnitude) control actions, which makes better adaptation pay off while it reduces overshoots that may destabilize the nonlinear controller.
@@ 7 Conclusion
In this work, we considered the use of deep polytopic autoencoders and state-dependent Riccati equations for the design of nonlinear controllers. We use the concept of linear parameter-varying approximations of nonlinear dynamical systems that are parametrized by deep polytopic autoencoders, and we show that by the particular choice of autoencoders, the parametrization is differentiable and has vanishing second-order terms. For the nonlinear feedback design, we extend the series expansion of state-dependent Riccati equations to second-order terms for which we provide computable expressions in form of indefinite Lyapunov and Riccati equations. Finally, we test the proposed framework on the classical control example modeling the flow past a cylinder.
The proposed nonlinear autoencoder with smooth clustering significantly improves the parametrization quality at very low dimensions. For controller design, we established that for second-order expansions of the state-dependent feedback law, only the first-order expansions of the reduced-order linear parameter-varying coefficient are relevant, which reduces the computational effort by large. Regarding the performance of the resulting feedback, the overall picture is diffuse. While the linear quadratic regulator approach, for suitable weight parameters, ensured stability also well away from the origin, the nonlinear feedback design based on different approximation schemes, clearly extended the basin of attraction. As the choice of the approximation was solely based on a single open-loop simulation in a time range that did not cover the full range of the closed-loop simulation, in future work, we expect to gain additional performance by optimizing the approximation in view of feedback control.
@@ Acknowledgments
Parts of this work have been carried out while Jan Heiland was with the Max Planck Institute for Complex Technical Systems, Magdeburg, and with the Otto von Guericke University Magdeburg.